OpenStudy (anonymous):

angle RTS a right angle and TU is an altitude if RU =8 and TU=4 what is the value of US

OpenStudy (anonymous):

US=2

OpenStudy (anonymous):

because...^RTU ~ ^ TUS RU/TU=TU/US US=TU^2/RU=16/8=2